OpenStudy (kayleigh98):

Given a triangle with sides measuring 18 inches, 21 inches, and 14 inches, find the measure of the angle with the greatest measure. A. 76 degrees B. 81 degrees C. 84 degrees D. 90 degrees

OpenStudy (eliesaab):

Use the formula \[ a^2=b^2 +c^2 - 2 bc \cos(A) \] Find \( \cos(A)\) then \(A\)

OpenStudy (eliesaab):

for a=21, b=14, c=18, you find \[ \cos(A)=\frac {79}{336} \]

OpenStudy (eliesaab):

Use a claculator to find that \[ A=76.4014 \]

OpenStudy (eliesaab):

The first choice is your answer